# Spontaneous phase locking in a broad-area semiconductor laser

### v1

@@ Abstract
Broad-area semiconductor lasers are employed in many high-power applications, however, their spatio-temporal dynamics is complex and intrinsically unstable due to the interaction of several transverse lasing modes. A dynamical and spatio-spectral analysis with ultra-high resolution of commercial broad-area lasers reveals multiplets of phase locked first- and second-order transverse modes that are spontaneously created by the nonlinear dynamics for a wide range of operation parameters. Phase locking between modes of different transverse order is demonstrated by comparing the linewidths of the lasing modes to that of their beat note as well as by a direct measurement of their phase fluctuation correlations. Since the laser and the setup lack any feature designed to induce locking and the overall dynamics is unstable, the observation of this spontaneous phase locking effect is very unexpected. Our findings indicate that chimera-like states featuring groups of synchronized oscillators coexisting with unsynchronized ones can spontaneously form in broad-area lasers, and may thus be found in a wider range of optical systems and beyond than previously assumed. Moreover, some of the phase locked modes do not even exist on the passive-cavity level, but are created by the nonlinear dynamics, an effect not previously observed in the context of chimera states.
@@ I Introduction
The dynamics of lasers is very rich, ranging from stable emission and periodic dynamics to deterministic chaos [1, 2, 3, 4]. Moreover, spatial pattern formation and complex spatio-temporal dynamics are observed for lasers featuring several transverse modes [5, 6], for example in gas lasers [7], fiber lasers [8, 9] and both surface-emitting [10, 11] and edge-emitting [12, 13, 14] semiconductor lasers. Such broad-area lasers (BALs) are needed to reach very high output powers beyond the limitations of single transverse-mode lasers. However, they are susceptible to spatio-temporal instabilities stemming from the interaction of different transverse lasing modes. While progress has been made in controlling spatio-temporal dynamics with special device designs [15, 16, 17, 18, 19, 20], the understanding or even just documentation of dynamical phenomena occurring in generic broad-area lasers remains far from complete.
@@
Beyond application-related considerations, BALs are hence interesting model systems for studying complex multimode dynamics. Important issues are the possibly chaotic nature of the spatio-temporal dynamics, how lasing modes organize in space, wavelength and time, and if synchronization or other forms of collective dynamics of lasing modes occur. An interesting form of collective dynamics are so-called chimera states [21, 22] in which groups of synchronized oscillators coexist with a large number of unsynchronized ones. Chimera states have been predicted numerically for arrays of coupled lasers [23] and demonstrated experimentally for semiconductor lasers with feedback [24, 25] and absorptive sections [26]. Similarly, synchronization of mode clusters was found for quantum dot [27] and quantum cascade [28] lasers. However, previous works concerned only lasers with a single transverse mode, whereas cluster synchronization of different transverse modes has not been studied before.
@@
We investigate a broad-area edge-emitting semiconductor laser with a quantum well as active medium. Semiconductor BALs suffer from filamentation, which is an intrinsic spatio-temporal instability caused by the combination of spatial-hole burning, carrier-induced index changes and diffraction [29, 12, 30, 31, 32, 13]. Filamentation is sometimes attributed to (partial) self-mode locking [30, 33, 34], and mode-locking between longitudinal and transverse modes can be induced by pulsed optical injection [33, 34]. However, phase- or mode-locking effects in normal, free-running quantum-well BALs have never been experimentally demonstrated even though they could explain the significant long-range correlations of the spatio-temporal dynamics [29, 12, 35, 36].
@@
We perform dynamical and spatio-spectral measurements with ultra-high resolution in order to identify the different transverse modes and find possible correlations between them. We observe a previously undocumented dynamical effect, the creation of multiplets of 1st- and 2nd-order transverse modes which are phase locked and that coexist with many unlocked modes. This chimera-like state is observed for two different lasers and in a wide range of operation parameters.
@@ II Experimental results
Figure 1: Emission properties of the BAL. a, RF-spectra of the total laser emission and b magnification around the particularly narrow frequency component near $2$ GHz. c, Spatio-spectral near-field image at $1000$ mA and its spatial and spectral projections (red). The small white bar indicates the free spectral range.
We investigate a commercial broad-area laser (see Appendix A.1) with continuous wave pumping in solitary operation, i.e., without optical injection, feedback or modulation. The lateral confinement in the $50~{}\mu$m wide and $1.5$ mm long cavity is due to a combination of gain and index guiding, and the laser threshold is $I_{th}=160$ mA.
Figure 1a shows the RF-spectra of the BAL as function of pump current. They were measured by coupling the complete laser beam into a multimode fiber connected to a fast photodetector (see Appendix A.2). The laser shows significant instabilities, and the RF-spectra feature a number of discrete and partially overlapping peaks on a broad background. The peaks typically have linewidths (FWHM) of $100$ MHz or more. However, one peak around $2$ GHz is significantly narrower (see magnification in Fig. 1b) with a linewidth below $10$ MHz for pump currents up to $600$ mA. It becomes broader and disappears as the pump current increases further (see Fig. 10). We surmise that the peaks in the RF-spectra are beat notes between different lasing modes [37], in which case the narrow peak could be explained by the beating of phase locked modes.
In order to confirm this, we performed spatio-spectral measurements using an imaging spectrometer with $30$ pm resolution (see Appendix A.2). The spatio-spectral image for $1000$ mA in Fig. 1c shows spatial patterns varying with the wavelength due to the presence of several transverse modes. The spectrally-integrated near-field intensity distribution has a top-hat profile with regular modulation which indicates the presence of high-order transverse modes (see Fig. 6 for more details). The spatially integrated spectrum shows only a single series of peaks with the longitudinal-mode spacing (free spectral range, FSR) of $25.9$ GHz, effectively hiding the presence of several transverse modes (see Fig. 5). More sophisticated grating-based setups enable spatio-spectral measurements with resolutions as good as $120$ MHz [38, 39, 40], which is sufficient to resolve the different transverse modes. However, in order to detect phase locking effects, we also need to measure the linewidths of the lasing modes which are of the order of $100$ MHz for multimode semiconductor lasers [41]. Hence, we adopt a heterodyning technique that permits even better spectral resolution [42, 43, 37] and measuring the phases of lasing modes [44].
@@
In order to perform spatially-resolved heterodyne measurements, we collect the BAL emission in the image plane of its output facet with a multimode fiber (MMF) whose lateral position is scanned (see Fig. 7). The signal of the BAL interferes with a single-mode reference laser on a fast photodetector to create beat notes from which we reconstruct the optical spectrum of the BAL (see Appendix B). The spectral resolution of the heterodyne measurement is fundamentally limited by the linewidth of the reference laser, which is $<1$ MHz.
@@
Figure 2c presents typical mode profiles, which show significant deviations from those expected in a passive Fabry-Perot (FP) cavity. First, the fundamental mode $q=1$ has a dip in the center and thus two peaks instead of one, in contrast to observations for other devices [39, 45, 38]. Second, the peak heights are not uniform. For example, the $q=3$ and $8$ modes have a higher peaks in the center, whereas the $q=7$ mode has an envelope with a dip in the center. Third, several modes have asymmetries even though the spectrally-integrated near-field intensity distributions are symmetric (see Fig. 6). In particular, the right (left) peak of the 1st (2nd) order mode is always higher. Thus, these two modes have complementary profiles as if they were competing for gain. It should be noted that the features discussed above are systematically found in all measurements. In a cavity with only index guiding in the lateral direction, the peak heights would of course be homogeneous [39], but an inhomogeneous current density as well as thermal lensing effects can significantly modify the mode profiles. In Ref. [39], thermal lensing was found to be the dominant mechanism, though no explanation for the observed asymmetries was given. Dynamical effects like spatial and spectral hole burning can play an important role as well.
@@
The spatially-resolved RF-spectrum in Fig. 2d shows that different frequency components exhibit different spatial patterns, which is expected if they result from beating of different transverse modes [37]. Figure 2e compares the spatially-averaged RF-spectrum (red line) to the beat notes (i.e., optical frequency differences) of the lasing modes (vertical gray lines) obtained from six heterodyne measurements at $1000$ mA. Almost all major peaks and frequency components in the RF-spectrum can thus be explained as beat notes. The widths of the peaks in the RF-spectrum (typically $100$ MHz and more) are also consistent with the widths of the lasing modes (see Fig. 2b). However, the significant offset of the RF-spectrum compared to the background signal (black line) over the whole frequency range as well as a few unexplained peaks indicate that dynamical effects beyond mode beating could contribute as well.
In conclusion, the experimental results for high pump currents are not unusual for a semiconductor BAL, even though some aspects of the RF-spectra and mode profiles merit further analysis. Now we turn our attention to lower pump currents to elucidate the creation of a narrow RF-peak by spontaneously phase-locked multiplets.
@@
Figure 3: Heterodyne measurement results at 350 mA. a, Reconstructed spatio-spectral image around a reference wavelength of $849.594$ nm. The green boxes mark a quintuplet and a triplet of 1st- and 2nd-order modes. b, Schematic of the quintuplet. The insets show the spatial profiles of the different modes, and the black arrows indicate their linewidths. The spacings and linewidth are given in MHz. c, Frequency of the narrow peak in the RF-spectrum (black) and NNN-spacings of multiplets (red) as function of the pump current. d, Linewidth of the narrow peak (black) and sum of linewidths of NNN-mode pairs. e, Linewidth reduction factor for different NNN-mode pairs.
@@
The spatio-spectral image for $350$ mA in Fig. 3a shows transverse modes up to order $q=6$. Surprisingly, we observe a quintuplet and a triplet of 1st- and 2nd-order modes (indicated by the green boxes), that is, additional lasing modes are created by the nonlinear dynamics. The schematic representation of the quintuplet in Fig. 3b shows two 1st-order modes followed by two 2nd-order modes and a fifth mode whose spatial profile does not allow a clear identification. Many such multiplets are found for $300$ to $650$ mA, always consisting of modes with $q=1$ and $2$, but also longitudinal mode groups with only one transverse mode of each order like for higher pump currents are observed. Most examples are triplets, though it is possible that sometimes weaker modes in the multiplets were not detected.
@@
Interestingly, the nearest-neighbor (NN) spacings indicated in Fig. 3b are not equal, instead, every second spacing is different. This means that four-wave mixing cannot be responsible for creating the multiplets since it would result in equidistant modes. More importantly, the next-nearest-neighbor (NNN) spacings are all the same and equal to the frequency of the narrow peak in the RF-spectrum highlighted in Fig. 1b. In other words, this peak results from to the beating of pairs of modes with different transverse order of which one was created by nonlinear laser dynamics. Figure 3c shows perfect agreement between the frequency of the peak in the RF-spectra (black line) and the NNN spacings (red) in the current range of interest.
@@
The linewidths of the lasing modes in the quintuplet are indicated by the double arrows in Fig. 3b. Every second mode has a clearly larger linewidth than its neighbors (see also Fig. 13). More importantly, the mode linewidths are significantly larger than that of the beat note, which is smaller than $10$ MHz in the regime up to $600$ mA (black symbols in Fig. 3d). If the phase fluctuations of the lasing modes were independent, the linewidth of their beat note would be equal to the sum of their linewidths (see Appendix D.1). Instead, the sum the the linewidths of the NNN-mode pairs (red symbols) is of the order of $100$ MHz in most cases. So the linewidth of the beat note is typically reduced by a factor (called linewidth reduction factor, LRF) ranging from $10$ to $50$ compared to the sum of the mode linewidths as shown in Fig. 3e. This means that the phase fluctuations of the NNN-mode pairs are highly correlated, i.e., these lasing modes are phase locked. This spontaneous phase locking is surprising since the BAL is in solitary operation without feedback, injection [33, 34], modulation or saturable absorbers [26]. It should be noted that the LRF shown in Fig. 3e increases till $600$ mA, i.e., the phase locking gets stronger, until it is lost around $650$ mA where the beat note becomes broader again (see also Fig. 10).
@@
The analysis of the linewidths is not a proof of phase locking of particular mode pairs since (i) it is not clear which NNN-mode pairs in the multiplets are phase locked or not and (ii) we cannot exclude the existence of mode pairs with narrow linewidths in other longitudinal mode groups that could explain the narrow beat note without phase locking. Furthermore we wonder if the NN-mode pairs may be weakly phase locked even though their beat note is not particularly narrow. Fortunately we can obtain direct proof of phase fluctuation correlations (PFCs) between different modes from the time-domain measurements of the heterodyne signals.
@@
Figure 4: Phase fluctuation correlations. a, Matrix of phase fluctuation correlations for the measurement shown in Fig. 3a. The green boxes indicate the quintuplet and the triplet. b, Phase fluctuation correlations as function of the joint signal-to-noise ratio for all mode pairs in data sets with multiplets. Pairs of first- and second-order modes are indicated in red, all other pairs in black. The gray rectangle indicates the range of not significant correlations of ±$\pm 0.1$. c, Phase fluctuation correlations of NN-mode pairs (green) and NNN-mode pairs (blue) only.
@@
The analysis to extract the phase fluctuations of lasing modes and calculate their correlations is explained in Appendix D. Figure 4(a) shows the PFC matrix for the lasing modes presented in Fig. 3, where the quintuplet and triplet are marked by green boxes. Most mode pairs exhibit negligibly small correlations, whereas significant correlations are observed between the modes of the multiplets. The highest correlations are found for NNN-mode pairs (second off-diagonal), all of which are clearly correlated. There are also correlations between the NN-mode pairs (first off-diagonal), but they are weaker. Moreover, it should be noted that some correlations are found between the modes of the quintuplet and the triplet which are one FSR away. These findings are exemplary for all the multiplets we found and are consistent with the linewidths of lasing modes and their beat notes as discussed above. They clearly prove that the modes created by the nonlinear dynamics are phase locked to their next-nearest neighbors, which suggests that a coherent effect is responsible for their creation.
@@
However, the measured PFCs are lower than could be expected from the linewidth reduction factors shown in Fig. 3e. For example, a LRF of $10$ corresponds to a PFC of $0.9$ (see Fig. 14), but such high values are not observed. In fact, the calculation of the PFCs is very sensitive to perturbations by measurement noise and interfering signals from the BAL dynamics. Figure 4b shows the PFCs of all measurements featuring multiplets as function of the joint signal-to-noise ratio (SNR), which is the mean of the SNRs of the two heterodyne signals (see Appendix D.3 for details). We see a clear increase of the maximal observed PFCs with the joint SNR. Hence, one should compare the PFC of a mode pair to that of other mode pairs with the same joint SNR in order to determine if the PFC can be considered high or not.
@@
Most mode pairs in Fig. 4 have PFCs in the range of ±$\pm 0.1$, which we consider as not significant and the involved modes as effectively uncorrelated. The comparison of correlations between modes of 1st and 2nd order (i.e., pairs with transverse quantum numbers $(1,1)$, $(1,2)$ or $(2,2)$, indicated in red) to all other mode pairs (indicated in black) shows that only the modes with $q=1$ and $2$ exhibit significant correlations, with a few exceptions to be discussed elsewhere. Figure 4c shows only the PFCs of NN- and NNN-mode pairs in multiplets, where the NNN-correlations are systematically higher than the NN-correlations for a given joint SNR. This confirms that the phase locking between NNN-modes is stronger compared to NN-pairs.

### v2

@@ Abstract
Broad-area semiconductor lasers are employed in many high-power applications, however, their spatio-temporal dynamics is complex and intrinsically unstable due to the interaction of several transverse lasing modes. A dynamical and spatio-spectral analysis with ultra-high resolution of commercial broad-area lasers reveals multiplets of phase-locked first- and second-order transverse modes that are spontaneously created by the nonlinear dynamics for a wide range of operation parameters. Phase locking between modes of different transverse order is demonstrated by comparing the linewidths of the lasing modes to that of their beat note and by a direct measurement of their phase fluctuation correlations. The spontaneous phase locking is unexpected since the overall dynamics is unstable and the system lacks any intentional feature to induce locking. This partially synchronized dynamical state with groups of synchronized and unsynchronized laser modes coexisting is similar to what is called a chimera state in networks of coupled oscillators, hence indicating that such states may exist in a wider range of systems than previously assumed. Moreover, some of the phase-locked modes do not exist on the passive-cavity level, but are created by the nonlinear dynamics, an effect not previously observed in the context of partially synchronized laser modes.
@@ I Introduction
The dynamics of lasers is very rich, ranging from stable emission and periodic dynamics to deterministic chaos [1, 2, 3, 4]. Moreover, spatial pattern formation and complex spatio-temporal dynamics are observed for lasers featuring several transverse modes [5, 6], including gas lasers [7], fiber lasers [8, 9] and both surface-emitting [10, 11] and edge-emitting [12, 13, 14] semiconductor lasers. Such broad-area lasers (BALs) are needed to reach high output powers beyond the limitations of single transverse-mode lasers. However, they are susceptible to spatio-temporal instabilities stemming from the interactions between transverse lasing modes mediated by the active medium. While progress has been made in controlling spatio-temporal dynamics with special device designs [15, 16, 17, 18, 19, 20], the documentation and understanding of dynamical phenomena occurring in simple broad-area lasers remains fragmentary.
@@
Beyond application-related considerations, BALs are hence interesting model systems for studying complex multimode dynamics. Important issues are the possibly chaotic nature of the spatio-temporal dynamics, how lasing modes organize in space, wavelength and time, and if synchronization or other forms of collective dynamics of lasing modes occur. An interesting form of collective dynamics exhibiting partial synchronization are chimera states [21, 22, 23] in which groups of synchronized oscillators coexist with unsynchronized ones. Chimera states have been predicted numerically for arrays of coupled lasers [24] and demonstrated experimentally for semiconductor lasers with feedback [25, 26] and absorptive sections [27]. Similarly, synchronization of mode clusters was found for a quantum cascade laser [28]. However, previous works concerned only lasers with a single transverse mode, whereas cluster synchronization of different transverse modes has not been studied before.
@@
We investigate a broad-area edge-emitting semiconductor laser with a quantum well as active medium. Semiconductor BALs suffer from intrinsic spatio-temporal instabilities caused by the interplay of spatial-hole burning, carrier-induced index changes and diffraction [29, 12, 30, 31, 32, 13]. Their spatio-temporal dynamics displays significant long-range correlations [29, 12, 33, 34] which has been attributed to partial self-mode locking [30, 35, 36]. While mode-locking between longitudinal and transverse modes can be induced by pulsed optical injection [35, 36], phase- or mode-locking effects in conventional, free-running quantum-well BALs have never been experimentally demonstrated.
@@
We perform dynamical and spatio-spectral measurements in order to identify the different transverse modes and find possible correlations between them. We observe an undocumented dynamical effect, the creation of multiplets of 1st- and 2nd-order phase-locked transverse modes that coexist with unlocked modes. This partially synchronized state is observed for two different lasers and in a wide range of operation parameters.
@@ II Experimental results
Figure 1: Emission properties of the BAL. a, RF-spectra of the total laser emission and b, magnification around the particularly narrow frequency component near $2$ GHz.
We investigate a commercial broad-area laser (see Appendix A.1) with continuous wave pumping in solitary operation (no optical injection, feedback or modulation). The lateral confinement in the $50~{}\mu$m wide and $1.5$ mm long cavity is due to a combination of gain and index guiding, and the laser threshold is $I_{th}=160$ mA (see Fig. 6).
Figure 1a shows the RF-spectra of the BAL as function of the pump current. They were measured by coupling the entire laser beam into a multimode fiber (MMF) connected to a fast photodetector (see Appendix A.3). The laser shows instabilities, and the RF-spectra feature a number of discrete and partially overlapping peaks on a broad background. Their typical linewidths (FWHM) are $100$ MHz or more. However, one peak around $2$ GHz is significantly narrower (see Fig. 1b) with a linewidth below $10$ MHz for pump currents up to $600$ mA. It becomes broader and disappears as the pump current increases further (see Fig. 13). We surmise that the peaks in the RF-spectra are beat notes between different lasing modes [37], in which case the narrow peak could be explained by the beating of phase-locked modes.
In order to confirm this, we perform spatio-spectral measurements, but the resolution of conventional imaging spectrometers is insufficient (see Appendix A.5). Sophisticated grating-based setups enable spatio-spectral measurements with resolutions as good as $120$ MHz [38, 39, 40, 41], which is sufficient to resolve the different transverse modes. However, in order to detect phase locking effects, we also need to measure the linewidths of the lasing modes which are of the order of $100$ MHz for multimode semiconductor lasers [42]. Hence, we adopt a heterodyning technique with even better spectral resolution [43, 44, 37] that additionally permits measuring the phases of lasing modes [45].
@@
For spatially-resolved heterodyne measurements, we collect the BAL emission in the image plane of its output facet with a MMF whose lateral position is scanned (see Fig. 10). The signal of the BAL interferes with a single-mode reference laser on a fast photodetector to create beat notes from which we reconstruct the optical spectrum of the BAL (see Appendix B). The spectral resolution of the heterodyne measurement is fundamentally limited by the linewidth of the reference laser ($<1$ MHz).
@@
Figure 2c presents typical mode profiles, which show significant deviations from those expected in a passive Fabry-Perot (FP) cavity with homogeneous refractive index distribution. First, the fundamental mode $q=1$ has a dip in the center and thus two peaks instead of one, in contrast to observations for other devices [39, 46, 38]. Second, the peak heights are not uniform. For example, the $q=3$ and $8$ modes have higher peaks in the center, whereas the $q=7$ mode has lower peaks in the center. Third, several modes have asymmetries even though the spectrally-integrated near-field intensity distributions are symmetric (see Fig. 7). In particular, the right (left) peak of the 1st (2nd) order mode is always higher. Thus, these two modes have complementary profiles as if they were competing for gain. It should be noted that the features discussed above are systematically found in all measurements. In a cavity with only index guiding in the lateral direction, the peak heights would of course be homogeneous [39], but an inhomogeneous current density as well as thermal lensing effects can significantly modify the mode profiles. In Ref. [39], thermal lensing was found to be the dominant mechanism, though no explanation for the observed asymmetries was given. Dynamical effects like spatial and spectral hole burning can play an important role as well.
@@
The spatially-resolved RF-spectrum in Fig. 2d shows that different frequency components exhibit different spatial patterns, which is expected for beating of different transverse modes [37]. Figure 2e compares the spatially-averaged RF-spectrum to the beat notes (i.e., optical frequency differences) of the lasing modes obtained from six heterodyne measurements at $1000$ mA. Almost all major peaks and frequency components in the RF-spectrum can thus be explained as beat notes. The widths of the peaks in the RF-spectrum are typically $100$ MHz and more and thus consistent with the widths of the lasing modes (see Fig. 2b). However, the significant offset of the RF-spectrum compared to the background signal over the whole frequency range and the few unexplained peaks indicate that dynamical effects beyond mode beating contribute as well.
@@
Figure 3: Heterodyne measurement results at 350 mA. a, Reconstructed spatio-spectral image around a reference wavelength of $849.594$ nm. The green boxes mark a quintuplet and a triplet of 1st- and 2nd-order modes. All modes are plotted with unit amplitude. b, Schematic of the quintuplet. The insets show the spatial profiles of the different modes, and the black arrows indicate their linewidths. The spacings and linewidths are given in MHz. c, Frequency of the narrow peak in the RF-spectrum (black) and NNN-spacings of multiplets (red) as function of the pump current. d, Linewidth of the narrow peak (black) and sum of linewidths of NNN-mode pairs (red). e, Linewidth reduction factor for different NNN-mode pairs, which indicates the local degree of synchronization.
@@
We next consider lower pump currents for which the narrow peak in the RF-spectra is observed. The spatio-spectral image for $350$ mA in Fig. 3a shows transverse modes up to order $q=6$. Surprisingly, we observe a quintuplet and a triplet of 1st- and 2nd-order modes (indicated by the green boxes), that is, additional lasing modes are created by the nonlinear dynamics. The schematic of the quintuplet in Fig. 3b shows two 1st-order modes followed by two 2nd-order modes and a fifth mode whose spatial profile does not allow a clear identification. Many such multiplets are found for $300$ to $650$ mA, always consisting of modes with $q=1$ and $2$, but also longitudinal mode groups with only one transverse mode of each order like for higher pump currents are observed.
@@
Interestingly, the nearest-neighbor (NN) spacings indicated in Fig. 3b are not equal, instead, every second spacing is different. This means that four-wave mixing cannot be responsible for creating the multiplets since it would result in equidistant modes. More importantly, the next-nearest-neighbor (NNN) spacings are all the same and equal to the frequency of the narrow peak in the RF-spectrum shown in Fig. 1b. So this peak results from to the beating of pairs of modes with different transverse order of which one was created by nonlinear dynamics. Figure 3c shows perfect agreement between the frequency of the narrow RF-peak and the NNN spacings in the current range of interest.
@@
The linewidths of the lasing modes in the quintuplet are indicated by the double arrows in Fig. 3b. Every second mode has a clearly larger linewidth than its neighbors (see also Fig. 16). More importantly, the mode linewidths are significantly larger than that of the beat note, which is smaller than $10$ MHz in the regime up to $600$ mA (see Fig. 3d). If the phase fluctuations of the lasing modes were independent, the linewidth of their beat note would be equal to the sum of their linewidths (see Section D.1). Instead, the sum of the linewidths of the NNN-mode pairs is of the order of $100$ MHz in most cases (Fig. 3d). So the linewidth of the beat note is typically reduced by a factor (called linewidth reduction factor, LRF) ranging from $10$ to $50$ compared to the sum of the mode linewidths as shown in Fig. 3e. This means that the phase fluctuations of the NNN-mode pairs are highly correlated, i.e., these lasing modes are phase-locked. This spontaneous phase locking is surprising since the BAL is in solitary operation without feedback, injection [35, 36], modulation or absorbing sections [27, 28]. It should be noted that the LRF shown in Fig. 3e increases till $600$ mA, i.e., the phase locking gets stronger, until it is lost around $650$ mA where the beat note becomes broader again (see also Fig. 13).
@@
The analysis of the linewidths is not a proof of phase locking of particular mode pairs since (i) it is not clear which NNN-mode pairs in the multiplets are phase-locked or not and (ii) we cannot exclude the existence of mode pairs with narrow linewidths in other longitudinal mode groups that could explain the narrow beat note without phase locking. Furthermore we wonder if the NN-mode pairs may be weakly phase-locked even though their beat note is not particularly narrow. Fortunately we can extract the phase fluctuation correlations (PFCs) between different modes from the time-domain measurements of the heterodyne signals.
@@
Figure 4: Phase fluctuation correlations. a, Matrix of phase fluctuation correlations for the measurement shown in Fig. 3a. The green boxes indicate the quintuplet and the triplet. b, Phase fluctuation correlations as function of the joint signal-to-noise ratio for all mode pairs in data sets with multiplets. Pairs of first- and second-order modes are indicated in red, all other pairs in black. The gray rectangle indicates the range of not significant correlations, $[-0.1,+0.1]$. c, Phase fluctuation correlations of NN-mode pairs (green) and NNN-mode pairs (blue) only.
@@
The calculation of the phase fluctuation correlations is explained in Appendix D. Figure 4(a) shows the PFC matrix for the lasing modes presented in Fig. 3, with the quintuplet and triplet marked by green boxes. Most mode pairs exhibit negligibly small correlations, whereas significant correlations are observed between the modes of the multiplets. The highest correlations are found for NNN-mode pairs (second off-diagonal), all of which are clearly correlated. There are also correlations between the NN-mode pairs (first off-diagonal), but they are weaker. Moreover, it should be noted that some correlations are found between the modes of the quintuplet and the triplet which are one FSR away. These findings are exemplary for all the multiplets and are consistent with the linewidths of lasing modes and their beat notes as discussed above. They clearly prove that the modes created by the nonlinear dynamics are phase-locked to their next-nearest neighbors, which suggests that a phase-sensitive nonlinear process is responsible for their creation.
@@
However, the measured PFCs are lower than expected from the linewidth reduction factors shown in Fig. 3e. For example, a LRF of $10$ corresponds to a PFC of $0.9$ (see Fig. 17), but such high values are not observed. In fact, the calculation of the PFCs is very sensitive to perturbations by measurement noise and interfering signals from the BAL dynamics. Figure 4b shows the PFCs of all measurements featuring multiplets as function of the joint signal-to-noise ratio (SNR), which is the mean of the SNRs of the two heterodyne signals (see Appendix D.3 for details). We see a clear increase of the maximal observed PFCs with the joint SNR. Hence, one should compare the PFC of a mode pair to that of other mode pairs with the same joint SNR in order to determine if the PFC can be considered high or not.
@@
Most mode pairs in Fig. 4 have PFCs in the range of $[-0.1,+0.1]$, which we consider as not significant. The comparison of correlations between modes of 1st and 2nd order [i.e., pairs with transverse quantum numbers $(1,1)$, $(1,2)$ or $(2,2)$] to all other mode pairs shows that only the modes with $q=1$ and $2$ exhibit significant correlations, with a few exceptions to be discussed elsewhere. Figure 4c shows only the PFCs of NN- and NNN-mode pairs in multiplets, where the NNN-correlations are systematically higher than the NN-correlations for a given joint SNR. This confirms that the phase locking between NNN-modes is stronger compared to NN-pairs.